OpenStudy (anonymous):

simplify the expression. the answer using exponents (5/3) to the power of 7

OpenStudy (unklerhaukus):

\[\left(\frac 5 3\right)^7=\frac{5^7}{3^7}=\frac{5\times5\times5\times5\times5\times5\times5}{3\times3\times3\times3\times3\times3\times3}\]

OpenStudy (anonymous):

hang on let me figure that out

OpenStudy (anonymous):

so the top is 78125?

OpenStudy (unklerhaukus):

it is

OpenStudy (anonymous):

2187 is he bottom

OpenStudy (anonymous):

that my final answer?

OpenStudy (unklerhaukus):

the bottom is 2187 yes.

OpenStudy (anonymous):

so what would my final answer be?

OpenStudy (unklerhaukus):

The question say to simplify, what is the simplest form ?

OpenStudy (anonymous):

ummm

OpenStudy (unklerhaukus):

i dont know what the questioner thinks is the most simple form it could be\[3^{-7}5^{7}\] i dont know